OpenStudy (anonymous):

What is the solution of log 2x – 3^125 = 3 ? x = 1 over 3 x = 1 x = 7 over 3 x = 4

OpenStudy (anonymous):

log basex a = b implies x^a = b

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

Is this correct? \[\Large \log (2x-3^{125}) = 3\]

OpenStudy (anonymous):

nah the base is 2x - 3

OpenStudy (anonymous):

so confuse >.<

OpenStudy (anonymous):

@telltoamit Only @gioanny39x can say that for sure. Gio, you need to confirm if I have the equation correct or not. Also, your formula is wrong. \[\Large log_B(X) = Y\] is the same as \[\Large X = B^{Y}\] Example: \[\Large log_10(1000) = 3\] B = 10 X = 1000 Y = 3 Using the 2nd form of the equation: \[\Large X = B^{Y}\] \[\Large 1000 = 10^{3}\]

OpenStudy (ash2326):

\[\log_{2x-3} 125=3\] Is this your question @gioanny39x ?

OpenStudy (anonymous):

only this gives the answer in options lol

OpenStudy (anonymous):